# Dark Energy Survey Year 6 Results: Cell-based Coadds and Metadetection Weak Lensing Shape Catalogue

### v1

@@ 8 Conclusion
Over the past decade, as the DES collaboration has worked diligently to improve our weak lensing analysis algorithms, our understanding of how to make weak lensing shear measurements has evolved significantly. During the SV analysis, one of the primary concerns in weak lensing measurements at the time was noise bias (e.g., Hirata & Seljak 2003; Kacprzak et al. 2012; Melchior & Viola 2012; Refregier et al. 2012). However, that analysis helped to reveal the importance of shear-dependent catalogue selection effects (e.g., Hirata & Seljak 2003; Bernstein & Jarvis 2002). In the analysis of Y1 and Y3 data, the collaboration moved to use weak lensing shear measurement techniques that correctly handled these noise bias and selection effects (Huff & Mandelbaum, 2017; Sheldon & Huff, 2017). The DES Y3 analysis introduced several additional innovations, including improved PSF modeling with Piff and extensive use of image simulations to calibrate cross-redshift, blending-induced shear calibration effects (Jarvis et al., 2021; Gatti, Sheldon et al. 2021; MacCrann, Becker et al. 2022). The DES Y3 analysis further helped to reveal the importance of shear-dependent detection effects and motivated the introduction of Metadetection itself (Sheldon et al., 2020). Finally, we have broadened our understanding of what a complete catalogue testing strategy entails and have made significant improvements to our catalogue testing strategies, including the introduction of more extensive statistical tests with PSF modeling residuals (Jarvis et al., 2016) and direct null tests on the cosmic shear two-point correlation functions (see, e.g., Becker et al., 2016; Amon et al., 2022).
@@ Appendix B Pre-PSF Gaussian (pgauss) Flux and Size Measurements
Our primary flux measure for each Metadetection detection is the pgauss flux measure from the ngmix package. This flux measure has the advantage of matching the effective flux apertures between different bands (up to PSF modeling errors), which is important for the measurement of object colors and photometric redshifts. This flux measure is not adaptive and so can be lower signal-to-noise than a flux measure with an aperture that is better matched to the object. We compute this flux measure, and a measure of object size, by closely following the procedure for the Fourier-space moments from Bernstein et al. (2016b) with some key modifications. This technique is also reminiscent of the Gaussian aperture (GaaP) fluxes from Kuijken (2008).
@@
[…]
@@ C.1 Junk Detections
    pgauss cut – As stated in Sec. 4.4, we removed any objects with $T^{\rm pgauss}>1.6-3.1\times T^{\rm pgauss}_{\rm err}$. The top panel of Fig. 21 shows $32\times 32$ pixel $gri$ false-color postage stamp images of randomly selected objects that are removed by this cut. The stamp is centered on the object we cut, which is not visible in many cases. Many of these objects are falsely detected due to background subtraction errors near bright foreground objects.
@@
Figure 21: Randomly selected objects that are removed from the catalogue by the “pgauss” cut (top) and the “super-spreader” cut (bottom). Each image is a $32\times 32$ pixel $gri$ false-color postage stamp around the centroid of the object we cut. In most cases, there is no obvious visible object at the center of the stamp.
@@ C.2 Stellar Contamination
Unfortunately, the estimates above are too low by a factor of several due to the fact that star-galaxy separation at faint magnitudes is poor in the Gold catalogue (Bechtol et al., 2025). Estimates from synthetic-source injection (Anbajagane, Tabbutt et al., 2025) indicate a higher fraction of stars around $\sim 0.74\%$. This fraction was estimated by computing the fraction of selected Metadetection objects from the injected sample that matched back to a star in the injection catalogue. However, this estimate is too high since the purity of the star-galaxy classifier for the injected catalogue at faint magnitudes is poor (Anbajagane, Tabbutt et al., 2025). Further, the image simulations also provide an estimate of the effects of stars on the catalogue. The estimates in Sec. 7 indicate no detectable residual additive bias and an upper bound on the change in the multiplicative bias $m$ of $<3.6\times 10^{-3}$ at three-sigma. Thus we conclude that the residual stellar contamination in the catalogue is small enough to be ignored.
@@ Appendix D PSF Leakage and Modeling Errors
We fit the contamination model to our data by first maximizing the likelihood of the two-point cross-correlations above with the Nelder-Mead algorithm (Nelder & Mead, 1965), and then running Markov Chain Monte Carlo (MCMC) with the emcee package (Foreman-Mackey et al., 2013) to obtain the posteriors. The 1D and 2D marginal posterior contours of the parameters are shown in Figure 22. The 2D contour includes $1,2,3\sigma$ contour lines, and the parameter values indicating no systematic contamination are shown as dotted lines. Most inferred parameters are consistent with zero PSF leakage and modeling errors, except $\beta^{(4)}$ and $\eta^{(24)}$. Each specific analysis using this shear catalogue will need to determine if the residual PSF modeling errors are significant enough to warrant further treatment via marginalizing over the contamination model parameters.
@@ 11 Kavli Institute for Cosmological Physics, University of Chicago, Chicago, IL 60637, USA
17 Laboratório Interinstitucional de e-Astronomia - LIneA, Rua Gal. José Cristino 77, Rio de Janeiro, RJ - 20921-400, Brazil

### v2

@@ 8 Conclusion
Over the past decade, as the DES collaboration has worked diligently to improve our weak lensing analysis algorithms, our understanding of how to make weak lensing shear measurements has evolved significantly. During the SV analysis, one of the primary concerns in weak lensing measurements at the time was noise bias (e.g., Hirata and Seljak 2003; Kacprzak et al. 2012; Melchior and Viola 2012; Refregier et al. 2012). However, that analysis helped to reveal the importance of shear-dependent catalogue selection effects (e.g., Hirata and Seljak 2003; Bernstein and Jarvis 2002). In the analysis of Y1 and Y3 data, the collaboration moved to use weak lensing shear measurement techniques that correctly handled these noise bias and selection effects (Huff and Mandelbaum, 2017; Sheldon and Huff, 2017). The DES Y3 analysis introduced several additional innovations, including improved PSF modeling with Piff and extensive use of image simulations to calibrate cross-redshift, blending-induced shear calibration effects (Jarvis, Bernstein, Amon, Davis, Léget, Bechtol, Harrison, Gatti, Roodman, Chang, Chen, Choi, Desai, Drlica-Wagner, Gruen, Gruendl, Hernandez, MacCrann, Meyers, Navarro-Alsina, Pandey, Plazas, Secco, Sheldon, Troxel, Vorperian, Wei, Zuntz, Abbott, Aguena, Allam, Avila, Bhargava, Bridle, Brooks, Carnero Rosell, Carrasco Kind, Carretero, Costanzi, da Costa, De Vicente, Diehl, Doel, Everett, Flaugher, Fosalba, Frieman, García-Bellido, Gaztanaga, Gerdes, Gutierrez, Hinton, Hollowood, Honscheid, James, Kent, Kuehn, Kuropatkin, Lahav, Maia, March, Marshall, Melchior, Menanteau, Miquel, Ogando, Paz-Chinchón, Rykoff, Sanchez, Scarpine, Schubnell, Serrano, Sevilla-Noarbe, Smith, Suchyta, Swanson, Tarle, Varga, Walker, Wester, Wilkinson, Wilkinson, and DES Collaboration 2021; Gatti, Sheldon, and others 2021; MacCrann, Becker et al. 2022). The DES Y3 analysis further helped to reveal the importance of shear-dependent detection effects and motivated the introduction of Metadetection itself (Sheldon et al., 2020). Finally, we have broadened our understanding of what a complete catalogue testing strategy entails and have made significant improvements to our catalogue testing strategies, including the introduction of more extensive statistical tests with PSF modeling residuals (Jarvis et al., 2016) and direct null tests on the cosmic shear two-point correlation functions (see, e.g., Becker et al., 2016; Amon et al., 2022).
@@ Appendix B Pre-PSF Gaussian (pgauss) Flux and Size Measurements
Our primary flux measure for each Metadetection detection is the pgauss flux measure from the ngmix package. This flux measure has the advantage of matching the effective flux apertures between different bands (up to PSF modeling errors), which is important for the measurement of object colors and photometric redshifts. This flux measure is not adaptive and so can be lower signal-to-noise than a flux measure with an aperture that is better matched to the object. We compute this flux measure, and a measure of object size, by closely following the procedure for the Fourier-space moments from Bernstein et al. (2016a) with some key modifications. This technique is also reminiscent of the Gaussian aperture (GaaP) fluxes from Kuijken (2008).
@@
    The integral of the product of two functions in real-space can be computed from their Fourier transforms: $\int dx^{2}G(x)W(x)=\int dk^{2}G(k)W(k)$. This relation is called the Plancherel theorem (Plancherel and Leffler, 1910).
@@
$M_{i}^{(n)}$ $=$ $\int dx^{2}x_{i}^{2}W(x)\left[I(x)\circledast P^{-1}(x)\right]$ (20)
$\propto$ $\int dk^{2}\frac{d^{n}W(k)}{dk_{i}^{n}}\left[I(k)/P(k)\right]$
@@ C.1 Junk Detections
    pgauss cut – As stated in Sec. 4.4, we removed any objects with $T^{\rm pgauss}>1.6-3.1\times T^{\rm pgauss}_{\rm err}$. The top panel of Fig. 21 shows $32\times 32$ pixel grigri false-color postage stamp images of randomly selected objects that are removed by this cut. The stamp is centered on the object we cut, which is not visible in many cases. Many of these objects are falsely detected due to background subtraction errors near bright foreground objects.
@@
Figure 21: Randomly selected objects that are removed from the catalogue by the “pgauss” cut (top) and the “super-spreader” cut (bottom). Each image is a $32\times 32$ pixel grigri false-color postage stamp around the centroid of the object we cut. In most cases, there is no obvious visible object at the center of the stamp.
@@ C.2 Stellar Contamination
Unfortunately, the estimates above are too low by a factor of several due to the fact that star-galaxy separation at faint magnitudes is poor in the Gold catalogue (Bechtol et al., 2025). Estimates from synthetic-source injection (Anbajagane, Tabbutt, Beas-Gonzalez, Yanny, Everett, Becker, Yamamoto, Legnani, De Vicente, Bechtol, Elvin-Poole, Bernstein, Choi, Gatti, Giannini, Gruendl, Jarvis, Lee, Mena-Fernández, Porredon, Rodriguez-Monroy, Rozo, Rykoff, Schutt, Sheldon, Troxel, Weaverdyck, Wetzell, Aguena, Alarcon, Allam, Amon, Andrade-Oliveira, Brooks, Rosell, Carretero, Chang, Crocce, da Costa, Pereira, Davis, Desai, Diehl, Dodelson, Doel, Drlica-Wagner, Ferté, Frieman, García-Bellido, Gaztanaga, Gruen, Gutierrez, Hartley, Herner, Hinton, Hollowood, Honscheid, Huterer, James, Krause, Kuehn, Lahav, Marshall, Miquel, Muir, Myles, Pieres, Malagón, Prat, Raveri, Samuroff, Sanchez, Cid, Sevilla-Noarbe, Smith, Suchyta, Tarle, Tucker, Walker, Wiseman, and Zhang, 2025) indicate a higher fraction of stars around $\sim 0.74\%$. This fraction was estimated by computing the fraction of selected Metadetection objects from the injected sample that matched back to a star in the injection catalogue. However, this estimate is too high since the purity of the star-galaxy classifier for the injected catalogue at faint magnitudes is poor (Anbajagane, Tabbutt, Beas-Gonzalez, Yanny, Everett, Becker, Yamamoto, Legnani, De Vicente, Bechtol, Elvin-Poole, Bernstein, Choi, Gatti, Giannini, Gruendl, Jarvis, Lee, Mena-Fernández, Porredon, Rodriguez-Monroy, Rozo, Rykoff, Schutt, Sheldon, Troxel, Weaverdyck, Wetzell, Aguena, Alarcon, Allam, Amon, Andrade-Oliveira, Brooks, Rosell, Carretero, Chang, Crocce, da Costa, Pereira, Davis, Desai, Diehl, Dodelson, Doel, Drlica-Wagner, Ferté, Frieman, García-Bellido, Gaztanaga, Gruen, Gutierrez, Hartley, Herner, Hinton, Hollowood, Honscheid, Huterer, James, Krause, Kuehn, Lahav, Marshall, Miquel, Muir, Myles, Pieres, Malagón, Prat, Raveri, Samuroff, Sanchez, Cid, Sevilla-Noarbe, Smith, Suchyta, Tarle, Tucker, Walker, Wiseman, and Zhang, 2025). Further, the image simulations also provide an estimate of the effects of stars on the catalogue. The estimates in Sec. 7 indicate no detectable residual additive bias and an upper bound on the change in the multiplicative bias $m$ of $3.6\times 10^{-3}$ at three-sigma. Thus we conclude that the residual stellar contamination in the catalogue is small enough to be ignored.
@@ Appendix D PSF Leakage and Modeling Errors
We fit the contamination model to our data by first maximizing the likelihood of the two-point cross-correlations above with the Nelder-Mead algorithm (Nelder and Mead, 1965), and then running Markov Chain Monte Carlo (MCMC) with the emcee package (Foreman-Mackey et al., 2013) to obtain the posteriors. The 1D and 2D marginal posterior contours of the parameters are shown in Figure 22. The 2D contour includes $1,2,3\sigma$ contour lines, and the parameter values indicating no systematic contamination are shown as dotted lines. Most inferred parameters are consistent with zero PSF leakage and modeling errors, except $\beta^{(4)}$ and $\eta^{(24)}$. Each specific analysis using this shear catalogue will need to determine if the residual PSF modeling errors are significant enough to warrant further treatment via marginalizing over the contamination model parameters.
@@ 11 Kavli Institute for Cosmological Physics, University of Chicago, Chicago, IL 60637, USA
17 Laboratório Interinstitucional de e-Astronomia - LIneA, Av. Pastor Martin Luther King Jr, 126 Del Castilho, Nova América Offices, Torre 3000/sala 817 CEP: 20765-000, Brazil
@@
19 Physik-Institut, University of Zürich, Winterthurerstrasse 190, CH-8057 Zürich, Switzerland
@@ 49 Max Planck Institute for Extraterrestrial Physics, Giessenbachstrasse, 85748 Garching, Germany
50 Department of Physics, University of Cincinnati, Cincinnati, Ohio 45221, USA